# Spin-Orbital Intertwined Topological Superconductivity in a Class of Correlated Noncentrosymmetric Materials

### v1

@@ Abstract
In this study, we propose an alternative route to achieving topological superconductivity (TSC). Our approach applies to a new class of correlated noncentrosymmetric materials that host two spin-split Fermi surfaces with identical spin textures due to a spin-orbital intertwined effect. Incorporating multi-orbital repulsive Hubbard interactions, we calculate the superconducting pairings of a minimal two-orbital effective model within a spin-fluctuation-mediated superconductivity framework. We find that, depending on the effective Rashba spin-orbit coupling (RSOC) strength and filling level, the Hubbard interaction can drive the leading pairing symmetry into the $A_{1}(S_{\pm})$, $B_{1}$, $B_{2}$ or $B_{2}(d_{\pm})$ irreducible representations (IRs) of the $C_{4v}$ point group. Notably, the $A_{1}(S_{\pm})$ pairing gives rise to a fully gapped TSC characterized by a $Z_{2}$ invariant, while the $B_{2}(d_{\pm})$ pairing results in a nodal TSC. Our analysis reveals that the fully gapped TSC is predominated by spin-singlet regardless of the presence of the spin-triplet components. This distinguishes our model from noncentrosymmetric materials with conventional Rashba-split band structures, where TSC typically emerges near the van Hove singularity and is primarily driven by $p$-wave or $f$-wave spin-triplet pairing. These features enhances its experimental accessibility, and we discuss potential experimental systems for its realization.
Introduction.— TSCs can host Majorana bound states, which are exotic quasi-particles obeying non-Abelian statistics. This unique feature makes TSCs fundamentally fascinating and potentially valuable for applications in quantum computing[1, 2, 3, 4, 5, 6, 7, 8, 9]. Over the past two decades, many efforts have been devoted to realizing TSCs in various material platforms[10, 11, 12, 13, 14, 15, 16, 17, 18]. The general strategy involves utilizing superconducting proximity effects in either momentum space or real space. The former, also known as connate TSCs, is typically achieved in multiband superconducting systems with topological boundary states, such as Fe(Se,Te), PbTaSe2 etc[13, 14, 15, 16, 17, 18, 19, 20]. The latter approach involves the artificial construction of topological insulator (or semiconductor)-superconductor heterostructures[10, 11, 12]. However, these material platforms pose experimental challenges, requiring high interface quality, a large induced superconducting gap, and precise control of the chemical potential.
@@
A comprehensive analysis of material systems exhibiting TSC reveals that their Fermi surfaces typically form robust spin textures, indicating explicit or potential topological characteristics in their electronic structures. Notably, materials such as CePt3Si, Y2C3, A2Cr3As3 (A=K, Rb, Cs)[24, 25, 26, 27, 28, 29], which lack inversion symmetry, display similar features due to RSOC. Comparable spin texture characteristics have also been experimentally observed in the Fermi surfaces of certain copper- and iron-based superconducting systems[30, 31]. These insights suggest that TSCs may emerge in correlated unconventional superconducting systems when RSOC is considered. However, current research on single-orbital Rashba-Hubbard models indicates that the small contribution of the spin-triplet component in these systems is typically insufficient to realize TSCs, except under specific conditions, such as fillings close to the van Hove singularity[32, 33, 34, 35].
To overcome the difficulties in realizing intrinsic TSCs, we propose an alternative approach by introducing a new degree of freedom, such as orbital, layer, or valley. For clarity, we focus on orbital degrees of freedom and consider a two-orbital Rashba-Hubbard model defined on a square lattice. Using the random phase approximation (RPA), we solve the superconducting problem of this model within the framework of spin-fluctuation-mediated superconductivity. We find that as the model parameters vary, the leading superconducting pairing can fall into the $A_{1}(S_{\pm})$, $B_{1}$, $B_{2}$ or $B_{2}(d_{\pm})$ of $C_{4v}$ point group. Interestingly, the $A_{1}(S_{\pm})$ pairing induces a sign-change gap function on the two Fermi surfaces with identical spin textures. This sign-change feature is primarily driven by the spin-singlet-pairing components regardless of the presence of the spin-triplet components. Meanwhile, the $A_{1}(S_{\pm})$ pairing leads to full-gap TSC states characterized by the topological invariant $Z_{2}$. This TSC is extraordinary and different from the ones driven by predominated $p$-wave or $f$-wave spin-triplet pairing. Additionally, the $B_{2}(d_{\pm})$ pairing results in a nodal TSC state with coexistence of zero-energy flat-band and dispersive Majorana states at the system’s boundary.
Figure 1: (a) The two-orbital ($d_{x^{2}-y^{2}},d_{xy}$) unconventional Rashba model defined on a square lattice with $C_{4v}$ symmetry. The relations between the parameters here and those in Eq. 2 are $t_{1/2}=(t_{x^{2}-y^{2}}\pm t_{xy})/2$ and $\lambda_{R}=(\lambda_{R1}+\lambda_{R2})/2$. (b)-(c) The representative band structure and Fermi surfaces for electron doping. The black arrows on the Fermi surfaces denote the effective spin and the color bar labels the renormalized orbital texture $\bra{FS}\tau^{z}\sigma^{0}\ket{FS}_{k}$/$\bra{FS}\tau^{z}\sigma^{0}\ket{FS}_{max}$. The parameters $t_{2}=0$, $\lambda_{R}=0.8$, $\epsilon=1$ and $\mu=-4.5$. (d-e) The calculated bare and RPA longitudinal spin susceptibilities $\chi_{0}^{zz}(\bm{q})$ and $\chi_{RPA}^{zz}(\bm{q})$. The transverse parts are weak[43]. The parameters $U=4.5$,$V=0.94U$ and $J=0.03U$. Other parameters are the same as those in (b) and (c).
Model and method.— The two-orbital unconventional Rashba-Hubbard model is given by
@@
$H_{0}=\sum_{k}\xi(k)-\lambda_{R}(\tau^{0}+\epsilon\tau^{3})[\bm{\sigma}\times\bm{\nabla_{k}}\gamma(k)]_{z}+\lambda\tau^{2}\sigma^{3}$ (2)
and
@@
$+J^{\prime}\sum_{i,\alpha<\beta}\sum_{\sigma}c^{\dagger}_{i\alpha\sigma}c^{\dagger}_{i\alpha\bar{\sigma}}c_{i\beta\bar{\sigma}}c_{i\beta\sigma}$
Here, $H_{0}$ can be defined on a lattice with the symmetry of $C_{3v}$, $C_{4v}$ or $C_{6v}$[36, 37, 38]. For clarity and simplicity, we consider square lattice with $C_{4v}$, as shown in Fig. 1 (a). The two orbital are fixed to be $d_{x^{2}-y^{2}}$ and $d_{xy}$. Then, $\xi(k)=2(t_{1}\tau^{0}+t_{2}\tau^{3})\gamma(k)-\mu$ with $\gamma(k)=\cos k_{x}+\cos k_{y}$ and $\mu$ denoting the chemical potential. $\lambda_{R}$ reperents the strength of effective RSOC, and a parameter $\epsilon$ is introduced to measure the anisotropy of two orbitals[39]. $\lambda$ is the strength of on-site SOC. To match the parameter settings in Fig. 1 (a), $t_{1/2}=(t_{x^{2}-y^{2}}\pm t_{xy})/2$ and $\lambda_{R}=(\lambda_{R1}+\lambda_{R2})/2$. In the following, $t_{1}$ is set to $-1$, and other parameters are in unit of $|t_{1}|$. $\lambda$ is set to 3 throughout the calculations. Pauli matrices $(\tau^{0},\bm{\tau})$ and $(\sigma^{0},\bm{\sigma})$ span in orbital and spin spaces, respectively. Then, $H_{0}$ spans in the bais $\Psi_{k}=(c_{x^{2}-y^{2},k,\uparrow},c_{x^{2}-y^{2}k,\downarrow},c_{xy,k,\uparrow},c_{xy,k,\downarrow})^{T}$ with $c_{\alpha k,\sigma}$ being the electron annihilation operator. $U$, $V$, $J$ and $J^{\prime}$ reperents the intra- and inter-orbital on-site repulsive Hubbard interactions, Hund’s coupling, and the pairing hopping term, respectively. $\hat{n}_{i\alpha\sigma}$ and $\bm{S}_{i\alpha}$ denote the spin-$\sigma$ electron density operator and spin operator for electrons in orbital $\alpha$ on $i$ site, respectively. For clarity, we adopt the conventional notations $J=J^{\prime}$ and $V=U-2J$ in the following discussion.
@@
According to the theory of fluctuation-mediated superconductivity[40, 41, 42], the effective static superconducting pairing interactions can be calculated by considering bubble and ladder diagrams[43], and are expressed as:
@@
Here, $\Delta_{l_{1}\sigma_{1}l_{4}\sigma_{4}}(\bm{k})$ is the superconducting pairing function and the leading superconducting instability can be identified by the largest positive eigen-value of $\lambda$.
@@
Fermionology and spin susceptibility.— Previous studies on the conventional Rashba-Hubbard model have focused on cases where the filling is near the van Hove singularity. This condition induces strong ferromagnetic fluctuations, which, in turn, favor the emergence of leading $p$- or $f$-wave spin-triplet superconducting states[32, 34, 35]. In this work, we relax this restriction and consider more general cases of electron doping beyond the van Hove singularity. Note that the hole doping is the mirror of the electron doping here. A typical band structure and Fermi surfaces under electron doping are shown in the Figs. 1 (b) and (c). Two Fermi surfaces have the same spin textures and apparent orbital textures, as shown in Fig. 1 (c).
The bare and RPA spin susceptibilities corresponding to the Fermi surfaces in Fig. 1 (c) can be calculated with Eqs. 4 and 6. The calculated results indicate that the transverse spin susceptibility $\chi^{+-}(\bm{q})$ is consistently much smaller than the longitudinal spin susceptibility $\chi^{zz}(\bm{q})$[43]. Therefore, we focus on $\chi^{zz}(\bm{q})$ as shown in Figs. 1 (d) and (e), which primarily determines the characteristics of spin fluctuations in the system. In both Figs. 1 (d) and (e), in addition to the four peaks near the $(\pi,\pi)$, eight new peaks emerge, as indicated by the white dashed arrows in these panels. These new peaks correspond to the nesting wave vectors of the two spin-split Fermi surfaces, which are consistent with the black dashed arrows shown in Fig. 1(a). This suggests that these eight new peaks play a crucial role in determining the characteristics of the potential superconducting state, distinguishing it from the single-orbital case. We also calculated the spin susceptibility for other doping cases[43] and found significantly different characteristics compared to the single orbital cases.

### v2

@@ Abstract
In this study, we propose an alternative route to realizing topological superconductivity (TSC). Our approach targets a new class of correlated noncentrosymmetric materials that host two spin-split Fermi surfaces due to spin-orbital intertwined effects. By investigating the superconducting pairings based on two-orbital Rashba-Hubbard model on a square lattice within a spin-fluctuation-mediated pairing framework, we find that, depending on model parameters, the leading superconducting state belongs to the $A_{1}(S_{\pm})$, $B_{2}$ or $B_{2}(d_{\pm})$ irreducible representations (IRs) of the $C_{4v}$ point group. Notably, the $A_{1}(S_{\pm})$ state features a sign-changed gap between the two spin-split Fermi surfaces and exhibits two key characteristics: (i) it is a parity-mixed state, with the ratio of even-parity to odd-parity components tunable via the Rashba spin-orbit coupling (RSOC) strength and onsite orbital-dependent potential; (ii) it is a fully gapped TSC, characterized by a $Z_{2}$ topological invariant. When RSOC is strong, the $B_{2}(d_{\pm})$ state can emerge and exhibits both parity-mixed and nodal TSC features. Further analysis reveals that the fully gapped TSC can be predominated by spin-singlet, despite of the presence of the spin-triplet components. This distinguishes our case from single-orbital noncentrosymmetric materials, where TSC arises from $p$-wave or $f$-wave spin-triplet pairing. In those systems, the Fermi level must be tuned near a van Hove singularity to enhance ferromagnetic fluctuations that drive the spin-triplet pairing. These distinctions enhance the experimental feasibility of our model.
I Introduction
TSCs can host Majorana bound states, which are exotic quasi-particles obeying non-Abelian statistics. This unique feature makes TSCs fundamentally fascinating and potentially valuable for applications in quantum computing[1, 2, 3, 4, 5, 6, 7, 8, 9]. Over the past two decades, many efforts have been devoted to realizing TSCs in various material platforms[10, 11, 12, 13, 14, 15, 16, 17, 18]. The general strategy involves utilizing superconducting proximity effects in either momentum space or real space. The former, also known as connate TSCs, is typically achieved in multiband superconducting systems with topological boundary states, such as Fe(Se,Te), PbTaSe2 etc[13, 14, 15, 16, 17, 18, 19, 20]. The latter approach involves the artificial construction of topological insulator (or semiconductor)-superconductor heterostructures[10, 11, 12]. However, these material platforms pose experimental challenges, requiring high interface quality, a large induced superconducting gap, and precise control of the chemical potential.
@@ I Introduction
A comprehensive analysis of material systems exhibiting TSC reveals that their Fermi surfaces typically form robust spin textures, indicating explicit or potential topological characteristics in their electronic structures. Notably, materials such as CePt3Si, Y2C3, A2Cr3As3 (A=K, Rb, Cs)[24, 25, 26, 27, 28, 29], which lack central inversion symmetry, display similar features due to RSOC. Comparable spin texture characteristics have also been experimentally observed in the Fermi surfaces of certain copper- and iron-based superconductors[30, 31]. These insights suggest that TSCs may emerge in correlated unconventional superconducting systems when RSOC is considered. However, current research on single-orbital Rashba-Hubbard models indicates that the small contribution of the spin-triplet component in these systems is typically insufficient to realize TSCs, except under specific conditions, such as fillings close to the van Hove singularity[32, 33, 34, 35].
To address the challenges in realizing intrinsic TSCs, we propose an alternative approach by introducing a new degree of freedom, such as orbital, layer, or valley. For clarity, we focus on orbital degrees of freedom and consider a two-orbital Rashba model, in which the two spin-split Fermi surfaces can exhibit either conventional or unconventional spin textures. This transition can be tuned by an on-site orbital-dependent potential together with onsite SOC, reflecting the evolution from the single-orbital limit to the two-orbital limit. Using the random phase approximation (RPA), we solve the superconducting problem of this model within the framework of spin-fluctuation-mediated superconductivity. We find that, depending on the strength of onsite SOC and effective RSOC, onsite orbital-dependent potential and filling level, the Hubbard interaction can drive the leading pairing symmetry into the $A_{1}(S_{\pm})$, $B_{2}$ or $B_{2}(d_{\pm})$ IRs of the $C_{4v}$ point group. Notably, the $A_{1}(S_{\pm})$ IR is conformed to host an $S_{\pm}$-wave pairing state, which lead to a fully gapped TSC characterized by a $Z_{2}$ topological invariant. Furthermore, the $S_{\pm}$-wave TSC is a parity-mixed state, with the parity-mixed ratio tunable via the strength of onsite SOC and effective RSOC and onsite orbital-dependent potential. This tunability offers a promising platform for exploring parity-mixed superconducting states. More significantly, the TSC can be predominatntly driven by spin-singlet component, despite of the spin-triplet contribution. We also find the $B_{2}(d_{\pm})$ state can emerge under the strong RSOC and exhibits both parity-mixed and nodal TSC features.
II Model and method
The two-orbital Rashba-Hubbard model[36, 37, 38, 39] is given by
@@ II Model and method
$H_{0}=\sum_{k}\xi(k)-\lambda_{R}(\tau^{0}+\epsilon\tau^{3})[\bm{\sigma}\times\bm{\nabla_{k}}\gamma_{1}(k)]_{z}+\lambda_{so}\tau^{2}\sigma^{3},$ (2)
in which, the SOC terms have the similar forms of those in the Kane-Mele model [40].RSOC is origined from $H_{RSOC}\propto(\nabla V\times p)\cdot\sigma$ under the condition $V(z)\neq V(-z)$ with $V(\pm z)$being the polarization fields along the ploar axis defined in Fig. 1 (a). Here, $H_{0}$ can be defined on a square lattice with $C_{4v}$ symmetry, as shown in Fig. 1 (a) . The two orbitals are fixed to be $d_{x^{2}-y^{2}}$ and $d_{xy}$. Then, $\xi(k)=2(t^{nn}_{1}\tau^{0}+t^{nn}_{2}\tau^{3})\gamma_{1}(k)+4(t^{nnn}_{1}\tau^{0}+t^{nnn}_{2}\tau^{3})\gamma_{2}(k)-\epsilon_{0}^{onsite}\tau^{3}-\mu$ with $\gamma_{1}(k)=\cos k_{x}+\cos k_{y}$, $\gamma_{2}(k)=\cos k_{x}\cos k_{y}$, $\epsilon_{0}^{onsite}$ indicating the difference of on-site orbital-dependent potential and $\mu$ denoting the chemical potential. $\lambda_{R}$ represents the strength of effective RSOC, and a parameter $\epsilon$ is introduced to measure the RSOC anisotropy of two orbitals[41]. $\lambda_{so}$ is the strength of on-site SOC. To match the parameter settings in Fig. 1 (a), $t^{nn/nnn}_{1/2}=(t^{nn/nnn}_{x^{2}-y^{2}}\pm t^{nn/nnn}_{xy})/2$ with $t^{nn/nnn}_{x^{2}-y^{2}/xy}$ denoting the nearest (nn) and next-nearest (nnn) neighbor hoppings for two $d_{x^{2}-y^{2}/xy}$ orbitals and $\lambda_{R}=(\lambda_{R1}+\lambda_{R2})/2$. In the following, $t^{nn}_{1}$ is set to 1-1, and other parameters are in unit of $|t^{nn}_{1}|$. $\lambda_{so}$ is set to $3+\epsilon_{0}^{onsite}$ throughout the calculations. Pauli matrices $(\tau^{0},\bm{\tau})$ and $(\sigma^{0},\bm{\sigma})$ span in orbital and spin spaces, respectively. Then, $H_{0}$ spans in the basis $\Psi_{k}=(c_{x^{2}-y^{2},k,\uparrow},c_{x^{2}-y^{2}k,\downarrow},c_{xy,k,\uparrow},c_{xy,k,\downarrow})^{T}$ with $c_{\alpha k,\sigma}$ being the electron annihilation operator.
The Hamiltonian $H_{0}$ in Eq. (2) yields the representative band structure, Fermi surfaces, and spin textures shown in Fig. 1. We find that when the parameter $\epsilon_{0}^{onsite}<-0.46$ or $\lambda_{so}<2.54$, the two spin-split Fermi surfaces display a conventional spin texture, indicating that the model reduces to the single-orbital limit. In contrast, when $\epsilon_{0}^{onsite}>-0.46$ or $\lambda_{so}>2.54$, the two spin-split Fermi surfaces exhibit an unconventional spin texture, signaling that the model is in the two-orbital limit. This property will impact the magnitude of transverse susceptibility and ultimately affect the effective inter-pocket interaction. The below discussions will cover both limits.
Figure 1: (a) The two-orbital ($d_{x^{2}-y^{2}},d_{xy}$) unconventional Rashba model defined on a square lattice with $C_{4v}$ symmetry. The polar axis is defined as the axis perpendicular to the square lattice plane, indicating that along the two opposite directions of this axis the system experiences different polarization fields $V(z)\neq V(-z)$. This asymmetry can in turn generate Rashba-type SOC. (b) The transition from single-orbital limit with conventional spin texture (left part in (b)) to two-orbital limit with unconventional spin texture (right part in (b)) controlled by $\epsilon_{0}^{onsite}$ or $\lambda_{so}$. The circles with cyan color denote the Fermi surfaces and the red arrows indicate the spin texture. (c) The band structures under different values of parameters $\epsilon_{0}^{onsite}$ or $\lambda_{so}$. From left to right, the parameters are set to be $\lambda_{so}=0$, $\lambda_{so}=1.5$ and $\lambda_{so}=3$ respectively. $\lambda_{so}=3+\epsilon_{0}^{onsite}$ is also fixed in order to prevent significant changes in the band structure. Other parameters are set as follows: $t^{nn}_{2}=-0.1$, $\lambda_{R}=0.5$, $\epsilon=1$, $t^{nnn}_{1}=0.0125$, $t^{nnn}_{2}=0.0375$, the occupied particle number per site is set to be $n=0.5$. Here, the coordinates of high-symmetry points are defined as follows: $\Gamma=(0,0)$; $X=(\pm\pi,0)$, $(0,\pm\pi)$; $M=(\pm\pi,\pm\pi)$.
The two-orbital Hubbard interactions can be expressed as:
@@
$+J^{\prime}\sum_{i,\alpha<\beta}\sum_{\sigma}c^{\dagger}_{i\alpha\sigma}c^{\dagger}_{i\alpha\bar{\sigma}}c_{i\beta\bar{\sigma}}c_{i\beta\sigma}.$
Here, $U$, $V$, $J$ and $J^{\prime}$ represent the intra- and inter-orbital on-site repulsive Hubbard interactions, Hund’s coupling, and the pairing hopping term, respectively. $\hat{n}_{i\alpha\sigma}$ and $\bm{S}_{i\alpha}$ denote the spin-$\sigma$ electron density operator and spin operator for electrons in orbital $\alpha$ on $i$ site, respectively. For clarity, we adopt the conventional notations $J=J^{\prime}$ and V=U2JV=U-2J in the following discussion.
[…]
@@
According to the theory of fluctuation-mediated superconductivity[42, 43, 44], the effective static superconducting pairing interactions can be calculated by considering bubble and ladder diagrams shown in Fig. A1, and are expressed as:
@@
Here, $\Delta_{l_{1}\sigma_{1}l_{4}\sigma_{4}}(\bm{k})$ is the superconducting pairing function and the leading superconducting instability can be identified by the largest positive eigen-value of $\lambda$. More details can be found in Appdenix A.
III Fermionology and spin susceptibility
@@
Previous studies on the conventional Rashba-Hubbard model have focused on cases where the filling is near the van Hove singularity. This condition induces strong ferromagnetic fluctuations, which, in turn, favor the emergence of leading $p$- or $f$-wave spin-triplet superconducting states[32, 34, 35]. In this work, we relax this restriction and consider more general cases of electron doping beyond the van Hove singularity. A typical band structure and Fermi surfaces under electron doping are shown in the Figs. 1 (c) and Figs. 2(a), (f) respectively.
To analyze the behavior of the spin susceptibilities, we define the scalar forms of longitudinal and transverse spin susceptibilities $\chi^{RPA}_{zz}(\bm{q})$ and $\chi^{RPA}_{+-}(\bm{q})$ as follows,
$\chi^{RPA}_{zz}(\bm{q})=\sum_{l_{1},l_{2}}\sum_{\{\sigma\}}\sigma^{z}_{\sigma_{1}\sigma_{2}}\sigma^{z}_{\sigma_{4}\sigma_{3}}\chi^{l_{1}\sigma_{1}l_{1}\sigma_{2}}_{l_{2}\sigma_{3}l_{2}\sigma_{4}}(\bm{q}),$ (12)
$\chi^{RPA}_{+-}(\bm{q})=\sum_{l_{1},l_{2}}\sum_{\{\sigma\}}\sigma^{+}_{\sigma_{1}\sigma_{2}}\sigma^{-}_{\sigma_{4}\sigma_{3}}\chi^{l_{1}\sigma_{1}l_{1}\sigma_{2}}_{l_{2}\sigma_{3}l_{2}\sigma_{4}}(\bm{q}).$
The RPA spin susceptibilities corresponding to the Fermi surfaces shown in Fig. 2 (a), (f), (k) and (p) can be calculated using Eqs. 4, 6 and 12. The difference between longitudinal and transverse susceptibility originates from the breaking of spin-rotation invariance induced by spin-splitting effects such as RSOC here. Specifically, in the single-orbital limit, the longitudinal susceptibility, as illustrated in Fig. 2(l), can be simplified as
$\chi^{RPA}_{zz}(\bm{q})=\sum_{\mu\nu,k\sigma,l_{1}l_{2}}C_{l_{1},l_{2}}^{\mu\nu}[1-\cos(\theta_{k}-\theta_{k+q})]\times g[E],$ (13)
where $e^{i\theta_{k}}=(\sin k_{y}+i\sin k_{x})/\sqrt{\sin^{2}k_{y}+\sin^{2}k_{x}}$, $g[E]=\frac{f(E_{\nu}(\bm{k}+\bm{q}))-f(E_{\mu}(\bm{k}))}{E_{\nu}(\bm{k}+\bm{q})-E_{\mu}(\bm{k})+i0^{+}}$ and the coefficient $C_{l_{1},l_{2}}^{\mu\nu}$ depend on $l_{1},l_{2},\mu,\nu$ and $k$ with $\mu$, $\nu$ denoting the band indices. For intra-pocket crossings, where k=(k+q)k=-(k+q), one has $\theta_{k}=\theta_{k+q}+\pi$, and therefore $1-\cos(\theta_{k}-\theta_{k+q})=2$, so that the susceptibility reaches its maximum. In contrast, for inter-pocket crossings this condition is not satisfied, and the susceptibility does not reach its maximum value. This explains why two distinct types of peaks are observed in Fig. 2(l). Meanwhile, the transverse susceptibility, shown in Fig. 2(m), can be written as
$\chi^{RPA}_{+-}(\bm{q})=\sum_{\mu\nu,k}|a^{\uparrow}_{\mu}(k)|^{2}|a^{\downarrow}_{\nu}(k+q)|^{2}\times g[E].$ (14)
The wave-function factor $|a^{\uparrow}_{\mu}(k)|^{2}|a^{\downarrow}_{\nu}(k+q)|^{2}$ is nearly constant (approximately 0.25) for both intra-pocket and inter-pocket crossings. As a result, three types of peaks appear, corresponding to intra-pocket crossing of the outer Fermi surface, inter-pocket crossing between different Fermi surfaces, and intra-pocket crossing of the inner Fermi surface.
[…]
We then examine the impact of transverse $\chi^{RPA}_{+-}(\bm{q})$. The numerical results indicate that for $\lambda_{so}=3$ (two-orbital limit) as shown in Fig. 2 (h) and (r), $\chi^{RPA}_{+-}(\bm{q})$ is approximately zero, leading to a comparable strength of effective interactions between intra-pocket and inter-pocket interactions, as shown in Fig. 2 (i) and (j) or (s) and (t). A stronger inter-pocket interaction favors a inter-pocket sign-change gap function making $\tau^{0}\sigma^{0}$ channel more likely to contribute to this sign change. Conversely, for $\lambda_{so}=0$ (one-orbital limit), the relatively strong $\chi^{RPA}_{+-}(\bm{q})$ depicted in Fig. 2 (c) and (m) renders the effective inter-pocket interaction being significantly weaker than the intra-pocket interaction as shown in Fig. 2 (d) and (e) or (n) and (o), thereby reducing the contribution of the $\tau^{0}\sigma^{0}$ channel to the inter-pocket sign-change behavior in the gap function. This indicates that the $\chi^{RPA}_{+-}(\bm{q})$ in Fig. 2 (c), (h), (m) and (r) plays a important role in influencing the contributions from different interaction channels, resulting in the $\tau^{0}\sigma^{0}$ channel of the two-orbital model being more significant and stable than that of the model in single-orbital limit, thereby distinguishing it from the latter.
IV Parity-mixed state and $S_{\pm}$-wave state